OpenStudy (anonymous):

Evaluate Sin^-1(sqrt(3)/2)

OpenStudy (anonymous):

find a number between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\) whose sine is \(\frac{\sqrt{3}}{2}\) it is a pretty common one